# Risk-Aware Distributional Intervention Policies for Language Models

### v1

@@ 5. Conclusion
Table 5. Ablation study: in the first scenario, we swap heads selected by RADIANT with ITI intervention, and vice versa; in the second scenario, we replace our risk-aware loss function with cross-entropy loss in training linear probe. Performed on TruthfulQA with Llama-7B.
Unintervened 21.15 22.16 25.58 40.54 2.13 0.00
ITI 26.52 28.03 27.78 43.59 2.20 0.07
1st scenario: Our linear probe + ITI intervention 26.88 28.00 29.00 44.00 2.17 0.04
1st scenario: ITI linear probe + our intervention 36.66 39.00 28.00 43.00 2.32 0.12
2nd scenario: Cross entropy loss 30.36 33.00 29.00 43.00 2.22 0.06
RADIANT 40.36 44.48 30.91 46.13 2.19 0.07
Appendix A Additional Experimental Details and Results
A.1. Analysis: The Effect of $\Gamma$ and $\alpha$ on the Performance of RADIANT
The hyperparameter $\alpha$ controls the conservativeness of the classifier in terms of the False Negative Rate. High values of $\alpha$ ensure that no undesirable content goes undetected. However, excessively large values of $\alpha$ may lead to trivial classifiers that classify all samples as undesirable. Such classifiers can be identified by checking if their False Positive Rate on the validation set is one. Therefore, for a given $\alpha$, along with other performance metrics, we report the average False Positive Rate and the average False Negative Rate across all trained classifiers on the validation set denoted $\overline{\text{FPR}}$ and $\overline{\text{FNR}}$.
In Table 7, we present metrics on the validation set while varying $\alpha$ within the set $\{1.0,1.5,2.0,2.5,3.0\}$. We use the base model Llama-7B. RADIANT’s performance improves as $\alpha$ increases until a significant drop occurs when trivial classifiers dominate at $\alpha=3.0$. This observation supports our approach of selecting $\alpha$ as high as possible without encountering the trivial-classifiers issue. However, the information score decreases as $\alpha$ increases. This decrease can be attributed to RADIANT becoming more conservative and avoiding providing uncertain information. In practice, depending on the information sensitivity of the application of LMs, we can select $\alpha$ as a trade-off between the accuracy of the information and the informativeness. For example, LMs in the medical or legal sectors should avoid providing incorrect or uncertain information, so high values of $\alpha$ are recommended.
We report the performance metrics of Llama-7B when varying $\Gamma$ in Table 6. This hyperparameter decides how much RADIANT post-intervention activations deviate from the original ones if detected as undesirable. We observe that the True score of RADIANT increases in $\Gamma$. This is because the increasing value of $\Gamma$ drives activations to reside deeper inside the desirable area, thus increasing the probability of desirable generation. However, the larger value of $\Gamma$ makes activations move farther from the original value, as shown by the increase in the CE and KL metrics. The extreme deviation from the original activations leads to inconsistency in semantics. It creates more non-natural sentences, which can be observed at $\Gamma=20$ with the drop in the Infomation score. Therefore, a reasonable score should balance between the True and Infomation scores.
In our implementation, for each pre-trained model, we perform a grid search where $\alpha$ ranges over $\{1.0,1.5,2.0,2.5\}$ and $\Gamma$ over $\{5,7.5,10,15,20\}$ to select the optimal combination based on the True * Info score in the validation set. After running RADIANT with various pretrained models, we find that the combination of $\Gamma=15$ and $\alpha=2.5$ performs effectively across most cases. Unless otherwise specified, we utilize these values for our experiments.
Table 6. The performance of RADIANT when varying $\Gamma$ and fixing $\alpha$ of 2.5.
@@ A.1. Analysis: The Effect of $\Gamma$ and $\alpha$ on the Performance of RADIANT
Table 7. The performance of RADIANT when varying $\alpha$ and fixing $\Gamma$ of 15.
$\alpha$ True * Info (%) $\uparrow$ True (%) $\uparrow$ Info (%) $\uparrow$ $\overline{\text{FPR}}$downarrow $\overline{\text{FNR}}$downarrow CE $\downarrow$ KL $\downarrow$
@@
A.2. The Transferability of Intervention Policies
We evaluated Llama-7B on NQOpen [22] using intervention vectors inherited from the TruthfulQA dataset. NQOpen contains approximately 3600 samples of question-answer pairs. Our intervention vectors show strong performance on out-of-distribution samples from the NQOpen dataset, shown in Table 8. This effectiveness is also observed with ITI, as noted in its original paper. Our experiment indicates that our intervention vectors offer superior transferability and generality compared to ITI’s. This experiment demonstrates the effectiveness of our method on larger datasets and highlights the generality of the computed intervention vectors for natural language tasks.
Table 8. Quantitative results of the transferability of RADIANT’s intervention on different datasets.
@@ A.2. The Transferability of Intervention Policies
A.3. The Effectiveness of RADIANT beyond the LLAMA Base Models
In this experiment, we study the performance of finetuning-free techniques, including ITI, RADIANT, and FSP, on Gemma-2B [40] and GPT-2 Large [35], which serve as alternative base models to the Llama model family. Table 9 shows that RADIANT using few-shot prompting outperforms other methods by a large gap. In particular, FSP + RADIANT improves the True * Info score of Gemma-2B and GPT-2 Large by 25.14% and 16.16%, respectively. Notably, FSP + RADIANT is superior to FSP + ITI in both True * Info and True and MC1 scores. Concurrently, RADIANT, implemented separately, outperforms ITI and FSP in terms of True * Info and True scores while only slightly behind in MC1 and MC2.
Table 9. Quantitative results of different intervention methods on TruthfulQA dataset, across different Language Models. Parameters of RADIANT: $\alpha=2.5,\Gamma=15$.
@@ A.3. The Effectiveness of RADIANT beyond the LLAMA Base Models
A.4. Toxicity mitigation task

### v2

@@ 4.2.1 Comparison between Finetuning-free Techniques
Table 2: Ablation study results: in the first scenario, we swap heads selected by RADIANT with ITI intervention, and vice versa; in the second scenario, we replace our risk-aware loss function with cross-entropy loss in training linear probe. Performed on TruthfulQA with Llama-7B.
Component Llama-7B Llama3-8B Llama2-chat-13B
Train the linear probe for one layer (s) 15.64 17.32 29.42
Compute intervention for one head (s) 52.33 58.43 55.67
Avg. increase in inference time per answer (%) 3.09 3.32 4.72
Table 3: Wall-clock time breakdown by components of RADIANT for different pretrained models.
4.2.3 Comparison against ITI with Representation Intervention Finetuning Methods.
We apply RADIANT and ITI on Llama-7B, Llama3-8B, and Llama2-chat-13B models, which were previously fine-tuned by LOFiT, a representation intervention finetuning method. The experimental results in the last big row of Table 1 show that RADIANT is better than ITI in improving correctness and informativeness across different Llama models. While ITI offers modest improvements in some instances, it generally lags behind RADIANT, especially in larger models.
4.3 Ablation Studies
We perform two ablation studies to demonstrate the effectiveness of our framework. Table 2 reports the performance of the Llama-7B + TruthfulQA dataset. In the first scenario, we select intervened heads using ITI, then compare our intervention approach versus ITI. We noticed that switching the head selection between RADIANT and ITI improved performance when the RADIANT intervention was applied, reaching 37% in the True * Info score. In the second scenario, the probing loss function is replaced by the popular binary cross-entropy loss. This scenario tests the impact of replacing the risk-aware loss function with cross-entropy loss, which resulted in moderate improvements but still fell short compared to RADIANT’s risk-aware loss in Section 2 (30.36% vs 40.36% in True*Info). Overall, these findings suggest that both the choice of intervention and the loss function play crucial roles in our framework.
4.4 Computational Cost
Our method is computationally cheap: for each head, our linear probes require one vector-vector multiplication, and our linear interventions require only one matrix-vector multiplication. To demonstrate this, we clocked the running time to calculate the intervention vectors on an A5000 GPU for the Llama-7B and Llama3-8B models and on two A5000 GPUs for Llama2-chat-13B and show the results in Table 3. Our intervention only slightly increases the running time of the inference process. In addition to its simplicity, the preprocessing of our framework for calculating intervention vectors is much less time-consuming and resource-intensive than fine-tuning methods.
4.5 Additional Experiments
We perform multiple additional benchmarks to demonstrate the effectiveness of RADIANT across diverse settings; however, due to space constraints, we have included the rest of these experiments in the Appendix. More specifically:
  • We conduct a hyperparameter analysis and compare our risk-aware loss function with weighted negative log-likelihood (Section A.2). We also report the results for the Kullback-Leibler (KL) and Cross-Entropy (CE) metric for generation (Section A.3). We find consistent improvements of RADIANT compared to other competitors.
  • We develop a parallel implementation (RADIANT-P) in Section A.4 that significantly reduces computational overhead while maintaining the performance of the intervention.
  • We conduct a transferability analysis (Section A.5) to demonstrate that RADIANT has better generalization to the NQOpen and TriviaQA datasets than ITI.
  • We test RADIANT on various model architectures, including Gemma and GPT models (Section B.1), Mistral and Qwen models (Section B.2), and sparse MoE architectures (Section B.3). We also compare RADIANT with supervised fine-tuning (Section B.4).
  • We evaluate RADIANT on other NLP tasks, including toxicity mitigation (Section C.1), long fact generation (Section C.2), and creative writing (Section C.3). RADIANT shows comparable performance to computationally intensive methods while maintaining better fluency.
5 Conclusion
We introduced RADIANT, an activation intervention method consisting of two components: (i) a layerwise probe to detect undesirable content and (ii) headwise interventions to rectify the head activations upon undesirably predicted outcomes. Unlike existing intervention methods, which can be scattered across different layers, our intervention is focused on a single layer of the network. This focus helps alleviate the distributional shifts of the activations in subsequent layers. Moreover, our headwise intervention aims to minimize perturbations to the activations while maintaining a reasonable guarantee of the intervention’s effectiveness. This is further demonstrated in empirical results, where our method outperforms the baseline intervention methods for various LMs.
Limitations and Social Impact. Our paper focuses on improving the truthfulness of LMs, and the results aim to improve trustworthy artificial intelligence. Apart from language generation, our paper can also be implemented in other domains for activation editing. However, it is important to acknowledge the potential misuse of our method. There exists a risk that adversarial actors could exploit our approach to transform truthful outputs into misleading or false information. This dual-use nature underscores the importance of ethical guidelines and safeguards in developing artificial intelligence. By promoting transparency and accountability in the use of our framework, we aim to raise awareness of the risks while maximizing the benefits of enhanced truthfulness in language generation.
Appendix A Additional Results on TruthfulQA Benchmark
A.1 Information about TruthfulQA Dataset
The TruthfulQA dataset is a Question-Answer dataset containing 817 questions that likely elicit false answers from humans due to common misconceptions. We follow the same data-processing used in Li et al. (2024b) and Yin et al. (2024) that splits the dataset into train/validation/test with the rate of $326/82/407$ questions and utilizes two-fold cross-validation. Each question has an average length of nine words and has two sets of desirable and undesirable answers. Following Li et al. (2024b), we separate the original dataset into 5918 question-answer pairs; each has a binary label, indicating desirability. Only pairs associated with questions in the training dataset are used to create our intervention policy, while those in the validation test are set aside for parameter tuning.
A.2 Additional Ablation Studies
A.2.1 Layer Selection Threshold with the Smooth Probing Loss
(b) FNR across layers for different value of regularization parameter $\alpha$ of the risk-aware loss in (2).
Figure 2: FNR and FPR metrics with different hyperparameters $\alpha$ across layers of Llama-7B.
A.2.2 Loss Function of the Classifier
To evaluate the effectiveness of our risk-aware loss function in (2), we conduct an ablation study that compares it with weighted negative log-likelihood (NLL) loss. The weighted NLL loss to train the classifier is defined as:
$\mathcal{L}_{\mathrm{NLL}}(\theta_{\ell h},\vartheta_{\ell h})$
$=-\sum_{i=1}^{N}[w_{1}y_{i}^{*}\log(p_{i})+w_{0}(1-y_{i}^{*})\log(1-p_{i})],$
where $p_{i}=\mathrm{sigmoid}(\vartheta_{\ell h}+\theta_{\ell h}^{\top}a_{\ell h,i})$ is the predicted probability for sample $i$, and $w_{1}$ and $w_{0}$ represent the weights applied to positive and negative samples, respectively. These weights are chosen to align with the class-imbalance handling and error-prioritization approach in our proposed loss. Specifically, we set
$w_{0}=\frac{1}{N_{0}},\quad w_{1}=\frac{\alpha}{N_{1}}$
to be consistent with the definitions of FPR and FNR in our formulation. We follow the same procedure to choose $\alpha$ as we do with RADIANT and run the experiment for three models: Llama-7B and Llama3-8B on the TruthfulQA dataset. The experimental setup is the same as in 4.1. The layers for intervention found by our framework with the NLL loss coincide with the layer found by our proposed loss (2) across all three models. We report the experimental results in Table 5.
Methods True*Info (%) $\uparrow$ True (%) $\uparrow$ MC1 $\uparrow$ MC2 $\uparrow$
Unintervened 21.15 22.16 25.58 40.54
RADIANT (NLL) 37.64 43.78 31.98 46.11
RADIANT 40.36 44.48 30.91 46.13
Table 4: Results for Llama-7B.
Methods True*Info (%) $\uparrow$ True (%) $\uparrow$ MC1 $\uparrow$ MC2 $\uparrow$
Unintervened 32.88 44.18 30.36 48.98
RADIANT (NLL) 38.36 47.25 32.12 49.83
RADIANT 37.78 50.82 33.82 52.98
Table 5: Results for Llama3-8B.
Our results show that, while the weighted NLL provides a strong baseline, RADIANT consistently outperforms it in most metrics and models. In particular, RADIANT improves performance in MC1 and MC2, demonstrating a stronger ability to guide model interventions. Although weighted NLL is a standard formulation, our findings suggest that RADIANT’s approach to error weighting provides additional benefits beyond simple likelihood-based weighting.
We provide theoretical intuition for why the RADIANT loss might lead to superior performance in practical classification tasks compared to weighted NLL in our problem. The weighted NLL measures the classification confidence via the negative logarithm of predicted probabilities. Thus, large mispredictions incur exponentially large penalties:
$\log(p_{i})\xrightarrow[p_{i}\to 0]{}\infty,\quad-\log(1-p_{i})\xrightarrow[p_{i}\to 1$$\infty$.
This exponential penalty will ensure that the boundary does not go deeper into the area of any classes because if that happens, many samples will have $p_{i}\to 1$, then the exponential penalty will pull it back. In our problem, the false negative samples are critical, so we expect the boundary to go deeper into the desirable area, and this loss is not suitable.
Methods Llama 7B Llama3-8B
RADIANT(NLL) 0.104 0.128
RADIANT 0.023 0.054
Table 6: False Negative Rate comparison between RADIANT loss and Weighted NLL loss denoted as RADIANT(NLL).
@@ A.2.2 Loss Function of the Classifier
Table 7: The performance of RADIANT when varying $\Gamma$ and fixing $\alpha$ of 2.5.
$\alpha$ True * Info (%) $\uparrow$ True (%) $\uparrow$ Info (%) $\uparrow$ $\overline{\text{FPR}}$ $\downarrow$ $\overline{\text{FNR}}$ $\downarrow$ CE $\downarrow$ KL $\downarrow$
@@
Table 8: The performance of RADIANT when varying $\alpha$ and fixing $\Gamma$ of 15.
Instead, RADIANT measures risk more directly (FPR, FNR), penalizing misclassification linearly in probability. Thus, while weighted NLL severely penalizes misprediction (logarithmic scale), RADIANT loss penalizes misclassification risk proportionally and is, therefore, more reasonable for our problem. Moreover, the exponential penalty potentially makes weighted NLL brittle or overly sensitive to outliers, which often exist in deep layers’ activations of complex architectures like transformers. This intuition is backed up by our empirical observation that the False Negative of the weighted NLL loss is consistently higher than our loss across two models, Llama 7B and Llama3-8B.
A.2.3 The Effect of $\Gamma$ and $\alpha$ on the Performance of RADIANT
The hyperparameter $\alpha$ controls the conservativeness of the classifier in terms of the False Negative Rate. High values of $\alpha$ ensure that no undesirable content goes undetected. However, excessively large values of $\alpha$ may lead to trivial classifiers that classify all samples as undesirable. Such classifiers can be identified by checking if their False Positive Rate in the validation set is one. Therefore, for a given $\alpha$, along with other performance metrics, we report the average False Positive Rate and the average False Negative Rate across all trained classifiers on the validation set denoted $\overline{\text{FPR}}$ and $\overline{\text{FNR}}$.
In Table 8, we present metrics on the validation set while varying $\alpha$ within the set $\{1.0,1.5,2.0,2.5,3.0\}$. We use the base model Llama-7B. RADIANT’s performance improves as $\alpha$ increases until a significant drop occurs when trivial classifiers dominate at $\alpha=3.0$. This observation supports our approach of selecting $\alpha$ as high as possible without encountering the trivial classifiers issue. However, the information score decreases as $\alpha$ increases. This decrease can be attributed to RADIANT becoming more conservative and avoiding providing uncertain information. In practice, depending on the information sensitivity of the application of LMs, we can select $\alpha$ as a trade-off between the accuracy of the information and the informativeness. For example, LMs in the medical or legal sectors should avoid providing incorrect or uncertain information, so high values of $\alpha$ are recommended.
We report the performance metrics of Llama-7B when varying $\Gamma$ in Table 7. This hyperparameter decides how much RADIANT post-intervention activations deviate from the original ones if detected as undesirable. We observe that the True score of RADIANT increases in $\Gamma$. This is because the increasing value of $\Gamma$ drives activations to reside deeper inside the desirable area, thus increasing the probability of desirable generation. However, the larger value of $\Gamma$ makes activations move farther from the original value, as shown by the increase in the CE and KL metrics. The extreme deviation from the original activations leads to inconsistency in semantics. It creates more non-natural sentences, which can be observed at $\Gamma=20$ with the drop in the Information score. Therefore, a reasonable score should balance the True and Information scores.
In our implementation, for each pre-trained model, we perform a grid search where $\alpha$ ranges over $\{1.0,1.5,2.0,2.5\}$ and $\Gamma$ over $\{5,7.5,10,15,20\}$ to select the optimal combination based on the True * Info score in the validation set. After running RADIANT with various pre-trained models, we find that the combination of $\Gamma=15$ and $\alpha=2.5$ performs effectively in most cases. Unless otherwise specified, we utilize these values for our experiments.
A.3 RADIANT Enhances Performance with Minimal Distribution Shift
Since headwise intervention applies a linear mapping to transform undesirable activations into desirable ones, it raises concerns about potential unintended shifts in meaning. To mitigate the risk of semantic drift, our intervention method is explicitly constrained via semidefinite programming to minimize the shift in activation space while enforcing a probabilistic guarantee that the modified activation crosses the classifier’s decision boundary.
To showcase that the meaning shift caused by RADIANT is minimal, we report two additional metrics: Kullback-Leibler (KL) divergence of the model’s next-token prediction distribution (pre- vs. post-intervention) and Cross-Entropy (CE) loss. These metrics quantify the shift in the generation distribution following the intervention. Lower values indicate smaller deviation from the original model’s behavior, reducing the likelihood of unnatural outputs or anomalous characters. The calculation details are provided in Li et al. (2024b). Due to space constraints, these metrics were omitted from the main paper.
We report the KL and CE values in Table 9, which includes all the settings reported in Table 1 from the main text. Our results show that RADIANT maintains comparable KL and CE values across various scenarios, demonstrating that it preserves the original distribution while significantly improving truthfulness.
Model Methods CE $\downarrow$ KL $\downarrow$
Llama-7B
Unintervened 2.13 0.00
ITI 2.20 0.07
FSP 2.13 0.00
NL-ITI 2.19 0.07
LITO 2.19 0.07
RADIANT (ours) 2.19 0.07
FSP + ITI 2.20 0.07
FSP + NL-ITI 2.20 0.07
FSP + LITO 2.20 0.07
FSP + RADIANT (ours) 2.20 0.08
Llama3-8B
Unintervened 2.38 0.00
ITI 2.50 0.13
FSP 2.38 0.00
NL-ITI 2.50 0.13
LITO 2.48 0.11
RADIANT (ours) 2.48 0.08
FSP + ITI 2.48 0.14
FSP + NL-ITI 2.49 0.14
FSP + LITO 2.54 0.17
FSP + RADIANT (ours) 2.52 0.15
Llama2-chat-13B
Unintervened 2.31 0.00
ITI 2.32 0.17
FSP 2.31 0.00
NL-ITI 2.33 0.17
LITO 2.34 0.18
RADIANT (ours) 2.35 0.18
FSP + ITI 2.33 0.13
FSP + NL-ITI 2.34 0.15
FSP + LITO 2.36 0.17
FSP + RADIANT (ours) 2.38 0.18
Alpaca
Base 2.81 0.00
+ ITI 2.88 0.14
+ RADIANT (ours) 2.81 0.13
Vicuna
Base 2.67 0.00
+ ITI 2.77 0.26
+ RADIANT (ours) 2.73 0.27
Llama variants + LOFiT
LOFiT (7B) 2.35 0.00
+ ITI 2.55 0.14
+ RADIANT (ours) 2.56 0.13
LOFiT (8B) 3.27 0.00
+ ITI 3.33 0.08
+ RADIANT (ours) 3.38 0.11
LOFiT (Chat-13B) 2.52 0.00
+ ITI 2.73 0.21
+ RADIANT (ours) 2.73 0.20
Table 9: Quantitative results of different intervention methods on the TruthfulQA dataset, across different Language Models and fine-tuning approaches. Parameters of RADIANT: $\alpha=2.5,\Gamma=15$.
A.4 Computational Cost – Paralled Version
This section studies the impact of ITI and RADIANT on the base models’ inference speed. From the theoretical aspect, it is evident that a head intervention of ITI, which is just a vector addition, is faster than that of RADIANT, which comprises a matrix multiplication and addition operator. This observation is proved again by the empirical results shown in Table 10. This table reports the average percentage increase in inference time per answer of ITI and RADIANT across the base models. It is observed that the normal version of RADIANT imposes more additional time in inference than ITI does. However, it should be noted that all RADIANT interventions are conducted on the same layer, while ITI interventions are carried out on multiple pairs of layer heads. This attribute of RADIANT allows us to parallel the interventions, which is impossible for ITI. We denote the parallel version of RADIANT as RADIANT-P and include it in Table 10. RADIANT-P offers the same decent results as RADIANT but imposes less computation cost to base models than RADIANT and ITI.
Base models ITI RADIANT RADIANT-P
Gemma-2B 2.53 6.82 1.75
GPT-2 Large 2.43 3.01 1.65
Llama-7B 2.46 3.09 1.45
Llama3-8B 2.51 3.32 1.55
Llama2-chat-13B 2.51 4.72 1.57
Table 10: The average percentage increase in inference time per answer of ITI and RADIANT across base models.
@@ A.4 Computational Cost – Paralled Version
Table 11: Quantitative results of the transferability of RADIANT’s intervention on different datasets.
A.5 Transferability Experiments
We train the intervention mappings for Llama-7B using the TruthfulQA dataset, but we evaluate its performance on the NQOpen dataset Kwiatkowski et al. (2019). The NQOpen dataset contains approximately 3600 question-answer pairs. Our intervention vectors show strong performance on the NQOpen dataset, shown in Table 11. This effectiveness is also observed with ITI, as noted in its original paper. Nevertheless, our experiment indicates that our intervention mappings offer superior transferability and generality compared to ITI’s. This experiment demonstrates RADIANT’s effectiveness and highlights the generality of the computed intervention for NLP tasks.
Appendix B Additional Comparison Results
B.1 Evaluation on GPT and Gemma Base Models
In this experiment, we study the performance of finetuning-free techniques, including ITI, RADIANT, and FSP, on Gemma-2B Team et al. (2024) and GPT-2 Large Radford et al. (2019), which serve as alternative base models to the Llama model family. Table 12 shows that RADIANT, using few-shot prompting, outperforms other methods by a large gap. In particular, FSP + RADIANT improves the True * Info score of Gemma-2B and GPT-2 Large by 25.14% and 16.16%, respectively. Notably, FSP + RADIANT is superior to FSP + ITI in both True * Info and True and MC1 scores. Concurrently, RADIANT, implemented separately, outperforms ITI and FSP in terms of True * Info and True scores, while only slightly behind in MC1 and MC2.
@@ B.1 Evaluation on GPT and Gemma Base Models
Table 12: Quantitative results of different intervention methods on TruthfulQA dataset, across different language models. Parameters of RADIANT: $\alpha=2.5,\Gamma=15$.
B.2 Evaluation on Mistral and Qwen Base Models
RADIANT has already been evaluated on base models beyond Llama, including Gemma-2B and GPT-2 Large (Appendix B.1). Because RADIANT relies solely on activation values (rather than logits or internal weights), it is directly applicable to any transformer-based model. To further validate this, we conducted an experiment on Mistral-7B-Instruct-v0.2 and Qwen2-7B-Instruct base models and evaluated the results using GPT4. The optimal intervention layers found by RADIANT are 18 and 25, respectively, for Qwen2-7B-Instruct and Mistral-7B-Instruct-v0.2.
Methods True * Info (%) $\uparrow$ True (%) $\uparrow$ MC1 $\uparrow$ MC2 $\uparrow$
Unintervened 80.05 67.20 55.69 70.93
ITI 76.87 62.30 54.23 70.41
RADIANT (ours) 81.94 68.17 55.23 71.75
Table 13: Experimental results of baselines for Mistral-7B-Instruct-v0.2.
Methods True * Info (%) $\uparrow$ True (%) $\uparrow$ MC1 $\uparrow$ MC2 $\uparrow$
Unintervened 64.13 50.55 40.02 59.89
ITI 73.68 50.30 40.88 60.86
RADIANT (ours) 77.51 53.05 40.39 61.17
Table 14: Experimental results of baselines for Qwen2-7B-Instruct.
Tables 13 and 14 show that RADIANT improves truthfulness and overall performance across different models. For Mistral-7B-Instruct-v0.2, RADIANT achieves the highest True * Info (81.94%) and True (68.17%) scores, outperforming both the baseline and ITI. For Qwen2-7B-Instruct, RADIANT leads in True * Info (77.51%) and True (53.05%).
Regarding comparisons with knowledge editing methods such as WISE Wang et al. (2024) and ConceptEdit Wang et al. (2023b), we emphasize that our work targets a different setting: inference-time intervention rather than parameter-based model editing. While model editing methods aim to alter stored knowledge via weight updates, RADIANT operates at inference time by activation intervention, enabling lightweight, reversible control over generation.
B.3 Evaluation on Sparse MoE Architectures
We show that RADIANT can be applied to a sparse Mixture of Experts. The core mechanism of RADIANT, layerwise and headwise activation editing, relies solely on access to activations at a given layer and does not depend on weight sharing or architectural uniformity. Moreover, since our method operates directly on attention head activations, and most sparse MoE designs apply sparsity to MLP layers rather than attention blocks, the architecture-level sparsity introduced by MoEs does not interfere with our intervention method. To verify our performance on this MoE architecture, we conduct experiments with a sparse MoE model, OLMoE-1B-7B-0924, from Muennighoff et al. (2024). We provide the results of our method, ITI, and base model for the TruthfulQA dataset as follows.
Table 15 shows that RADIANT works well for the OLMoE-1B-7B-0924 model, boosting the True score by nearly 8%. RADIANT outperforms ITI regarding True * Info, True, MC1, and MC2. This showcases its effectiveness in generating truthful answers in architectures like Sparse MoEs.
Unintervened 23.43 31.42 28.92 45.42 2.67 0.00
ITI 25.52 32.74 28.48 45.50 2.84 0.13
RADIANT (ours) 28.56 39.17 30.13 47.05 2.68 0.12
Table 15: Performance on TruthfulQA with OLMoE-1B-7B-0924.
B.4 Comparison with Supervised Fine-Tuning
Supervised fine-tuning (SFT) attempts to align LLMs with human preferences (Ouyang et al., 2022b). Given a prompt, SFT encourages the model to generate desirable answers and reduce the likelihood of generating undesirable answers by optimizing the cross-entropy loss. However, SFT’s requirement to fine-tune all LLM parameters demands substantial GPU resources for the back-propagation operations. Due to computational constraints, we can only perform SFT on the GPT2-large, the smallest model in our experiments.
Parameter Value
learning_rate 0.00002
weight_decay 0
adam_beta1 0.8
adam_beta2 0.999
adam_epsilon $1\times 10^{-8}$
max_grad_norm 1
batch_size 32
epochs_num 5
lr_scheduler_type linear
Table 16: Parameter values for SFT.
Table 12 highlights the advantages of inference-time methods like ours: by avoiding gradient computation or backpropagation, they offer a lightweight, fast, versatile, and economical way to improve the performance of LLMs. This is especially useful in low-resource scenarios. Because Llama-7B is used as a base model for many of our experiments, we also include the results of SFT on Llama-7B for comparison, but it is worth noting that the number is taken from the ITI paper Li et al. (2024b). Since our evaluation framework differs from ITI in terms of the GPT-judge and GPT-info models, which is attributed to the fact that these models in the ITI paper are no longer available on OpenAI, the results may not be fair for comparison. From Table B.1, SFT achieves the best performance in terms of MC metrics and reaches a high score of True * Info and True. Regarding the True score, RADIANT still outperforms SFT in the individual and integrating versions with FSP, offering 38.73% and 40.41% correct answers, respectively. When combined with FSP, RADIANT achieves 35.36% in True * Info score, surpassing SFT but requiring fewer resources. For the implementation of SFT, we use the SFTTrainer framework [^2] https://huggingface.co/docs/trl/en/sft_trainer, one of the most popular frameworks for this algorithm. While we remained almost the default parameters proposed by the library, we had to tune many important parameters like learning rate, parameters of the Adam optimizer, weight decay, and so on, to get a consistent and stable fine-tuned model. Some important parameters for SFT are reported in Table 16, while its best performance is shown previously in Table 12. This observation strongly supports the practicability of RADIANT, which only necessitates tuning two key hyperparameters $\alpha$ in the probe loss (2), and $\Gamma=\Phi^{-1}(1-\gamma)$ in the computation of the intervention map (4). A detailed analysis of these parameters to provide insight into their impact is presented in Appendix A.2.3. This section offers useful insights and guidelines for selecting values for any new models. Furthermore, compared to other methods like ITI, the grid search on two hyperparameters like ours is efficient and reasonable, so it is not harder to tune the hyperparameters of RADIANT than other previous works.
Appendix C Experiments on Other NLP Tasks
C.1 Toxicity Mitigation Benchmark